# Exploration vs Exploitation: Rethinking RLVR through Clipping, Entropy, and Spurious Reward

### v2

@@ 1 Introduction
The recent emergence of Large AI Reasoning Models (e.g., Kimi-K2, OpenAI-o1, and DeepSeek-R1 (Team-2025-Kimi; Jaech-2024-OpenAI; Guo-2025-Deepseek)) has been driven by reinforcement learning with verifiable rewards (RLVR) (Li-2025-Review). In RLVR, a verifier compares the model’s rollout against a deterministic ground-truth solution, especially in mathematics and other STEM domains, providing outcome rewards. This verifiability has enabled models to achieve competitive and human-level performance on challenging benchmarks, such as the International Mathematical Olympiad (Huang-2025-Gemini). Among RLVR methods, Group Relative Policy Optimization (GRPO) (Shao-2024-Deepseekmath) has become particularly popular due to its computational simplicity and memory efficiency.
@@
These properties give RLVR a distinctive exploration–exploitation regime. In classical RL, spurious rewards, which are misaligned with the true outcome reward (e.g., random noise), would be expected to hinder exploitation by injecting randomness that encourages suboptimal actions. Yet in RLVR, they have been observed to improve performance in Qwen-Math models (Shao-2025-Spurious), a phenomenon attributed to upper-clipping bias that disproportionately amplifies high-prior responses, consistent with contamination effects reported on MATH500 (Wu-2025-Memorization). Conversely, entropy minimization, which reduces policy entropy to yield more deterministic, high-confidence rollouts, has been widely adopted in RLVR and empirically linked to consistent gains (Zhang-2025-EMPO; Zhao-2025-Learning; Cui-2025-Entropy; Fu-2025-How). Notably, Agarwal-2025-Entropy and Gao-2025-One directly optimize entropy as an objective and report substantial improvements even without verifiable feedback. These findings point to an RLVR-specific paradox: discouraging exploitation through spurious rewards and discouraging exploration through entropy minimization can both enhance validation accuracy, underscoring learning dynamics that depart from classical RL intuitions.
@@ 2.1 Group Relative Policy Optimization
RLVR assigns a binary outcome-based reward $\mathbf{r}(\mathbf{x},\mathbf{y})$ to a sampled response $\mathbf{y}$ from prompt $\mathbf{x}$ by comparing it against the ground-truth answer $\mathbf{y}^{\star}$. To learn an optimized policy via these reward, policy gradient methods (Williams-1992-Simple; Sutton-1998-Introduction) aim to maximize
@@
In this context, GRPO (Shao-2024-Deepseekmath) and its variants (Yu-2025-DAPO; Liu-2025-Understanding; Chu-2025-GPG; Zhang-2025-GVPO; Chen-2025-Stepwise; Li-2025-Knapsack) estimate policy gradients using groups of samples. For each prompt $\mathbf{x}$, GRPO draws a set $\{\mathbf{y}^{(i)}\}_{i=1}^{G}$ from $\pi_{\theta_{\textnormal{old}}}$. We denote $\mathbf{y}^{(i)}_{t}$ as the $t$-th token of $i$-th sample $\mathbf{y}^{(i)}$ and $\mathbf{h}_{t}^{(i)}:=\{\mathbf{x},\mathbf{y}_{<t}^{(i)}\}$ and optimize the clipped objective as follows,
@@
Following (Cui-2025-Entropy), we use the softmax policy update framework, and one typical iteration amounts to one-step exponentiation update with $G$ rollouts $\{\mathbf{y}^{(i)}\}_{i=1}^{G}$ as follows,
@@ 2.2 Spurious Reward for RLVR
We examine several empirical findings related to random rewards. Notably, Shao-2025-Spurious reports striking performance gains on MATH500 for the Qwen-Math family when models are fine-tuned using the random reward defined in Definition 2.3. However, similarly large improvements are not observed for several other model families. Wu-2025-Memorization likewise find substantial contamination in Qwen-Math on the MATH500 validation benchmark, hypothesizing that the apparent gains under random reward largely stem from reinforcing memorized or contaminated trajectories. In particular, Shao-2025-Spurious attributes these gains to the PPO-style upper-clipping bias, formalized as follows,
@@
However, Oertell-2025-Heuristics challenge this interpretation, arguing that the reported gains arise from algorithmic heuristics and evaluation artifacts; in their experiments, random-reward fine-tuning does not consistently improve reasoning and can even degrade it. These conflicting findings highlight how little is currently understood about RLVR learning dynamics and motivate two central questions: (i) Can random rewards improve model performance, and under what conditions? (ii) Does clipping bias provide a meaningful learning signal, and if not, what role does it actually play? Following prior work, our empirical analysis also focuses primarily on MATH500. We further discuss the broader implications of random-reward training for general reinforcement learning settings in Appendix A.
@@ 2.3 LLM Policy Entropy
Policy entropy ${\mathcal{H}}(\pi_{\theta})$ quantifies the diversity of a policy’s action distribution. A high-entropy policy allocates probability more evenly across actions, producing a wider variety of sampled responses, whereas a low-entropy policy concentrates probability on a small subset of actions, resulting in more deterministic behavior (Li-2025-Preserving).
@@
Recent works in RLVR has begun to examine how policy entropy influences model performance. A common perspective emphasizes avoiding “entropy collapse” to prevent premature convergence to a low-diversity, suboptimal policy (Yu-2025-DAPO). At the token level, Wang-2025-Beyond similarly highlight the importance of minority high-entropy tokens for effective reasoning. Yet several studies report the opposite pattern: reducing entropy can be beneficial. Agarwal-2025-Entropy explicitly optimize an entropy-minimization objective and observe performance improvements, and Cui-2025-Entropy even propose a monotonic relationship in which lower entropy yields better performance. These conflicting findings raise a second fundamental question: (iii) Is there a direct causal relationship between policy entropy and policy performance?
Beyond empirical observations, Cui-2025-Entropy provide a theoretical analysis by deriving the following estimate of the one-step change in policy entropy:
@@ 3.2 Model-specific Evaluation
Following the hyperparameter configuration of Shao-2025-Spurious, we train Qwen2.5-Math-7B on the DeepScaleR dataset (Luo-2025-Deepscaler) using random rewards drawn from $\textnormal{Bernoulli}(\frac{1}{2})$. The training setup uses a batch size of $128$, group size of $16$, decoding temperature $1.0$, clipping ratio $0.2$, learning rate $5\times 10^{-7}$, and KL coefficient 0.
We run multiple consecutive experiments with and without clipping using the verl framework (Sheng-2025-Hybridflow). The resulting training trajectories on the MATH500 validation set, together with the clipping activation fraction over training, are shown in Figure 1. We adopt the default training prompt from verl, which instructs the model to enclose its final answer in a box for verifier validation (see Appendix A for further discussion). Notably, for Qwen2.5-Math-7B, the clipping activation rate is substantially lower than what is typically observed in other base models:
@@ 4 Clipping and Policy Entropy
We provide two theoretical results describing policy entropy under unclipped and clipped training (§4.1). As discussed in §2.3, the approximation in Eq.˜5 from Cui-2025-Entropy becomes inaccurate when clipping or random rewards are present. Our analysis incorporates both clipping and initial policy skewness, yielding a more precise characterization of entropy dynamics. In §4.2, we validate these results through extensive experiments and targeted case studies. In §4.3, we interpret clipping as a mechanism that implicitly reduces entropy and caution – supported by empirical evidence – against conflating entropy reduction with improved performance.
@@ 4.1 One-step Policy Entropy Change under Random Rewards
We first present the unclipped-training dynamics in Theorem˜4.1, with detailed statement and proof in Section˜C.3, where we also identify the conditions that permit entropy growth.
@@
Next, we analyze training dynamics with upper clipping in Theorem˜4.3; detailed statements, proof and entropy decay verification under practical parameters are deferred to Section˜C.3.
@@
where $c_{G}$, $\Phi$ and $R(\eta)$ are defined as the same as in Theorem˜4.1; for other constants, see Section˜C.3.
@@
As shown in Figure 2 (Middle), our experiments confirm that policy entropy consistently decreases over time under random rewards. In contrast, disabling clipping leads to entropy increasing during training (Figure 2, Left). Existing approaches to counter early-stage entropy collapse rely on regularization techniques that merely slow the decay (Wang-2025-Harnessing; Yao-2025-Diversity; Zheng-2025-First; Cheng-2025-Reasoning). Our finding that one can actively increase policy entropy – while also improving validation performance – suggests a complementary strategy: using spurious-reward setups to more effectively preserve and modulate entropy. This highlights a promising direction for combining true and spurious rewards to better balance exploration and exploitation in RLVR.
@@ 5 Reward Misalignment: Who can Benefit from Random Rewards?
From empirical observations in this and prior work, we note two consistent patterns under random-reward training. First, in line with Shao-2025-Spurious, weaker models tend to improve less—and importantly, model strength is dataset-dependent: a model that performs well on an easier benchmark may struggle on a harder one. Second, as baseline accuracy increases (e.g., approaching 70%), training dynamics become noticeably smoother, whereas models starting around 50% accuracy exhibit substantially more oscillation. To explain when and why a model may improve under random rewards, we analyze the phenomenon through the lens of reward misalignment. As a warm-up, we introduce a simple probabilistic model that captures this mechanism in the binary outcome-reward (ORM) setting, converting the observed behavior into a tractable misalignment analysis.
@@
As reported by Shao-2025-Spurious, base Llama models reliably degrade during random-reward training across trials. Under the reward-misalignment perspective, stronger models should benefit more and are thus more likely to improve. We test this by evaluating a stronger distilled Llama variant, whose base and teacher models both exhibit contamination on MATH500. As shown in Figure 4 (Middle), using a rollout length of $8192$ tokens and matching all other hyperparameters to the Qwen-Math configuration, we observe improvements comparable to those in Figure 2. In contrast, the weaker and potentially contaminated Qwen-Math model (Figure 4, Left) fails to achieve similar gains. These results indicate that validation-set contamination does not account for the improvements under random rewards, nor is the effect specific to Qwen-Math. Figure 4 (Right) summarizes the percentage improvements across the model results in Figure 1 (Left) and Figure 4 (Left and Middle).
@@ Appendix A Related Works
We provide broader context on how prior work in reinforcement learning for classical settings has leveraged spurious rewards to facilitate training. Spurious reward signals are closely linked to challenges in generalization (Zhang-2018-Dissection; Koch-2021-Objective; Langosco-2022-Goal). While these works illustrate deliberate uses of such signals, spurious rewards may also arise unintentionally, leading to reward misspecification and reward hacking (Pan-2022-Effects); similar reward hacking has been documented in Tien-2023-Causal. A second relevant thread traces back to potential-based reward shaping (PBRS) (Ng-1999-Invariance), which introduces additional or misaligned rewards in principled ways to preserve the optimality of desired behaviors. More recently, Random Network Distillation (RND) (Burda-2019-Exploration) emerged as a leading exploration mechanism, with subsequent extensions such as Ma-2025-Perturbation. In this same context, numerous other works propose reward signals (often spurious with respect to the true task objective) that encourage an agent to explore the state space in ways that eventually uncover genuine rewards (Pathak-2017-Curiosity; Zhang-2021-NovelD; Wang-2023-LIBERTY; Li-2024-Breaking). Spurious rewards have thus played a substantial role in improving exploration. One prominent theoretical foundation is posterior sampling (Russo-2014-Learning), which motivates exploration through uncertainty and has been generalized to broader RL settings (Chen-2024-SICNN; Xu-2025-RLQESA).
@@
We now turn to recent works that study spurious rewards in RLVR. Although these works report broadly similar empirical phenomena, their experimental configurations differ in important ways. In Shao-2025-Spurious, the prompt omits the standard Qwen-style instruction to place the final answer in a box. As they note, Qwen-Math is highly sensitive to prompt formatting, and such differences can substantially shift baseline performance. In contrast, our experiments follow the default Qwen prompt used in verl (Sheng-2025-Hybridflow), which explicitly instructs the model to place the final answer in a boxed expression—mirroring the RLVR verifier in verl, which extracts the boxed answer for scoring and reward assignment.
Apart from this prompt choice, we match all other hyperparameters in Shao-2025-Spurious. Oertell-2025-Heuristics, however, adopt a markedly different configuration: (i) a rollout-length cap of 1024 tokens (well below the 4096-token context window of Qwen-Math), (ii) a different training dataset (MATH (Hendrycks-2021-Measuring) instead of DeepScaleR (Luo-2025-Deepscaler)), (iii) a significantly smaller learning rate ($1\times 10^{-7}$ versus $5\times 10^{-7}$ in Shao-2025-Spurious), and (iv) a reduced batch size (64 versus 128). The smaller learning rate changes the effective update magnitude, and the smaller batch size yields noisier estimates of the stochastic reward distribution. Given these differences, the empirical results reported across prior works are not directly comparable.
@@
We further comment on potential contamination in the Qwen2.5-Math models. As reported by Wu-2025-Memorization, contamination in Qwen-Math has been observed only on validation sets (e.g., MATH500). Beyond these findings, the official Qwen2.5-Math technical report (Qwen2.5-Math, Table 1) also acknowledges possible contamination arising from the close similarity between the training and validation sets of the MATH dataset (Hendrycks-2021-Measuring). Their training corpus comprises two components: (i) CoT data synthesis, which includes GSM8K, MATH, and NuminaMath (Qwen2.5-Math, §3.1.1), and (ii) a tool-integrated reasoning dataset containing GSM8K, MATH, CollegeMath, NuminaMath, MuggleMath, and DotaMath (Qwen2.5-Math, §3.1.2). In contrast, our experiments employ the DeepScaleR training set, which consists exclusively of selected questions from AMC, AIME, Omni-Math, and Still (Luo-2025-Deepscaler). None of these datasets appear in the training sources listed for Qwen2.5-Math. Therefore, we believe that our training data does not overlap with the datasets used to train Qwen2.5-Math and is thus not contaminated.
@@
Agarwal-2025-Entropy demonstrate that token-level entropy minimization can substantially improve LLM reasoning without verifiable feedback, arguing that reduced entropy increases model confidence and reveals latent reasoning capability. This mechanism parallels clipped training under random rewards, where updates primarily modulate entropy rather than exploit informative rewards. However, we show that entropy minimization alone may drive the policy toward low-entropy yet suboptimal solutions; hence, entropy should be viewed as a stabilizing regularizer rather than a replacement for genuine RLVR signals.
Related work explores entropy through the lens of self-confidence. In particular, Prabhudesai-2025-Maximizing use low-entropy rollouts as implicit rewards, achieving gains across diverse benchmarks, while Gao-2025-One show that even a single unlabeled example can improve reasoning via entropy reduction. Methods such as EMPO (Zhang-2025-EMPO) and Zhao-2025-Learning similarly enhance performance in unsupervised settings by amplifying model confidence. Niekerk-2025-Post further construct preference datasets from confidence scores, achieving RLHF-level improvements without human feedback. In this context, Cui-2025-Entropy propose a simple but influential empirical relationship between policy entropy $\mathcal{H}$ and model performance $R$, fit across extensive experiments:
@@
To mitigate early-stage entropy collapse, several works propose alternative strategies. Shen-2025-On analyze why entropy regularization suffers from limited benefit in RLVR training for LLMs by attributing it to the vast response space and the sparsity of optimal outputs, and then introduce an adaptive entropy-control method using a clamped entropy bonus with automatically tuned coefficients. Song-2025-Outcome show that ORM yields induces sharp reductions in output entropy and diversity (as shown by lower pass@n scores), and propose outcome-level entropy bonuses to counteract it. Prior works (Wang-2025-Harnessing; Yao-2025-Diversity; Zheng-2025-First; Cheng-2025-Reasoning) also develop additional techniques for controlling entropy during RLVR training.
@@
Proximal policy optimization (PPO) (Schulman-2017-Proximal) has become standard for reward-based policy updates in LLM training and remains a core component of RLHF. However, since PPO requires loading and maintaining four separate models during training, it is computationally and memory intensive. This has motivated the development of lighter-weight and adapted policy-gradient updates (Li-2024-Remax; Ahmadian-2024-Back; Shao-2024-Deepseekmath; Guo-2025-Deepseek). In parallel, advances in verifiable reward construction (Cobbe-2021-Training; Uesato-2022-Solving; Zelikman-2022-STaR; Singh-2023-Progprompt; Hosseini-2024-Vstar; Lightman-2024-Verify; Wang-2024-Math; Luo-2024-Improve; Setlur-2025-Rewarding; Zhang-2025-Lessons) have enabled reinforcement learning to substantially improve LLM reasoning, particularly in mathematical problem solving. Beyond training algorithms, recent work also explores post-processing and collaborative strategies to strengthen reasoning performance. Kay-2025-Consensus and Zhao-2025-Majority propose consensus-based and answer-aggregation methods within multi-model frameworks. Chen-2025-SelfQuestioning introduce a self-questioning paradigm for iterative refinement, while Park-2025-MAPoRL develop an online multi-agent collaborative reinforcement learning framework.

### v3

@@ 1 Introduction
The recent emergence of Large AI Reasoning Models (e.g., Kimi-K2, OpenAI-o1, and DeepSeek-R1 (Kimi, 2025; Jaech et al., 2024; Guo et al., 2025)) has been driven by reinforcement learning with verifiable rewards (RLVR) (Li et al., 2025c). In RLVR, a verifier compares the model’s rollout against a deterministic ground-truth solution, especially in mathematics and other STEM domains, providing outcome rewards. This verifiability has enabled models to achieve competitive and human-level performance on challenging benchmarks, such as the International Mathematical Olympiad (Huang and Yang, 2025). Among RLVR methods, Group Relative Policy Optimization (GRPO) (Shao et al., 2024) has become particularly popular due to its computational simplicity and memory efficiency.
@@
These properties give RLVR a distinctive exploration–exploitation regime. In classical RL, spurious rewards, which are misaligned with the true outcome reward (e.g., random noise), would be expected to hinder exploitation by injecting randomness that encourages suboptimal actions. Yet in RLVR, they have been observed to improve performance in Qwen-Math models (Shao et al., 2025), a phenomenon attributed to upper-clipping bias that disproportionately amplifies high-prior responses, consistent with contamination effects reported on MATH500 (Wu et al., 2025). Conversely, entropy minimization, which reduces policy entropy to yield more deterministic, high-confidence rollouts, has been widely adopted in RLVR and empirically linked to consistent gains (Zhang et al., 2025b; Zhao et al., 2025b; Cui et al., 2025; Fu et al., 2025). Notably, Agarwal et al. (2025) and Gao et al. (2025) directly optimize entropy as an objective and report substantial improvements even without verifiable feedback. These findings point to an RLVR-specific paradox: discouraging exploitation through spurious rewards and discouraging exploration through entropy minimization can both enhance validation accuracy, underscoring learning dynamics that depart from classical RL intuitions.
@@ 2.1 Group Relative Policy Optimization
RLVR assigns a binary outcome-based reward $\mathbf{r}(\mathbf{x},\mathbf{y})$ to a sampled response $\mathbf{y}$ from prompt $\mathbf{x}$ by comparing it against the ground-truth answer $\mathbf{y}^{\star}$. To learn an optimized policy via these reward, policy gradient methods (Williams, 1992; Sutton and Barto, 1998) aim to maximize
@@
In this context, GRPO (Shao et al., 2024) and its variants (Yu et al., 2025; Liu et al., 2025b; Chu et al., 2025; Zhang et al., 2025a; Chen et al., 2025c; Li et al., 2025b) estimate policy gradients using groups of samples. For each prompt $\mathbf{x}$, GRPO draws a set $\{\mathbf{y}^{(i)}\}_{i=1}^{G}$ from $\pi_{\theta_{\textnormal{old}}}$. We denote $\mathbf{y}^{(i)}_{t}$ as the $t$-th token of $i$-th sample $\mathbf{y}^{(i)}$ and $\mathbf{h}_{t}^{(i)}:=\{\mathbf{x},\mathbf{y}_{<t}^{(i)}\}$ and optimize the clipped objective as follows,
@@
Following Cui et al. (2025), we use the softmax policy update framework, and one typical iteration amounts to one-step exponentiation update with $G$ rollouts $\{\mathbf{y}^{(i)}\}_{i=1}^{G}$ as follows,
@@ 2.2 Spurious Reward for RLVR
We examine several empirical findings related to random rewards. Notably, Shao et al. (2025) reports striking performance gains on MATH500 for the Qwen-Math family when models are fine-tuned using the random reward defined in Definition 2.3. However, similarly large improvements are not observed for several other model families. Wu et al. (2025) likewise find substantial contamination in Qwen-Math on the MATH500 validation benchmark, hypothesizing that the apparent gains under random reward largely stem from reinforcing memorized or contaminated trajectories. In particular, Shao et al. (2025) attributes these gains to the PPO-style upper-clipping bias, formalized as follows,
@@
However, Oertell et al. (2025) challenge this interpretation, arguing that the reported gains arise from algorithmic heuristics and evaluation artifacts; in their experiments, random-reward fine-tuning does not consistently improve reasoning and can even degrade it. These conflicting findings highlight how little is currently understood about RLVR learning dynamics and motivate two central questions: (i) Can random rewards improve model performance, and under what conditions? (ii) Does clipping bias provide a meaningful learning signal, and if not, what role does it actually play? Following prior work, our empirical analysis also focuses primarily on MATH500. We further discuss the broader implications of random-reward training for general reinforcement learning settings in Appendix A.
@@ 2.3 LLM Policy Entropy
Policy entropy ${\mathcal{H}}(\pi_{\theta})$ quantifies the diversity of a policy’s action distribution. A high-entropy policy allocates probability more evenly across actions, producing a wider variety of sampled responses, whereas a low-entropy policy concentrates probability on a small subset of actions, resulting in more deterministic behavior (Li et al., 2025a).
@@
Recent works in RLVR has begun to examine how policy entropy influences model performance. A common perspective emphasizes avoiding “entropy collapse” to prevent premature convergence to a low-diversity, suboptimal policy (Yu et al., 2025). At the token level, Wang et al. (2025b) similarly highlight the importance of minority high-entropy tokens for effective reasoning. Yet several studies report the opposite pattern: reducing entropy can be beneficial. Agarwal et al. (2025) explicitly optimize an entropy-minimization objective and observe performance improvements, and Cui et al. (2025) even propose a monotonic relationship in which lower entropy yields better performance. These conflicting findings raise a second fundamental question: (iii) Is there a direct causal relationship between policy entropy and policy performance?
Beyond empirical observations, Cui et al. (2025) provide a theoretical analysis by deriving the following estimate of the one-step change in policy entropy:
@@ 3.2 Model-specific Evaluation
Following the hyperparameter configuration of Shao et al. (2025), we train Qwen2.5-Math-7B on the DeepScaleR dataset (Luo et al., 2025) using random rewards drawn from $\textnormal{Bernoulli}(\frac{1}{2})$. The training setup uses a batch size of $128$, group size of $16$, decoding temperature $1.0$, clipping ratio $0.2$, learning rate $5\times 10^{-7}$, and KL coefficient 0.
We run multiple consecutive experiments with and without clipping using the verl framework (Sheng et al., 2025). The resulting training trajectories on the MATH500 validation set, together with the clipping activation fraction over training, are shown in Figure 1. We adopt the default training prompt from verl, which instructs the model to enclose its final answer in a box for verifier validation (see Appendix A for further discussion). Notably, for Qwen2.5-Math-7B, the clipping activation rate is substantially lower than what is typically observed in other base models:
@@ 4 Clipping and Policy Entropy
We provide two theoretical results describing policy entropy under unclipped and clipped training (§4.1). As discussed in §2.3, the approximation in Eq.˜5 from Cui et al. (2025) becomes inaccurate when clipping or random rewards are present. Our analysis incorporates both clipping and initial policy skewness, yielding a more precise characterization of entropy dynamics. In §4.2, we validate these results through extensive experiments and targeted case studies. In §4.3, we interpret clipping as a mechanism that implicitly reduces entropy and caution – supported by empirical evidence – against conflating entropy reduction with improved performance.
@@ 4.1 One-step Policy Entropy Change under Random Rewards
We first present the unclipped-training dynamics in Theorem˜4.1, with detailed statement and proof in Appendix C.3, where we also identify the conditions that permit entropy growth.
@@
Next, we analyze training dynamics with upper clipping in Theorem˜4.3; detailed statements, proof and entropy decay verification under practical parameters are deferred to Appendix C.3.
@@
where $c_{G}$, $\Phi$ and $R(\eta)$ are defined as the same as in Theorem˜4.1; for other constants, see Appendix C.3.
@@
As shown in Figure 2 (Middle), our experiments confirm that policy entropy consistently decreases over time under random rewards. In contrast, disabling clipping leads to entropy increasing during training (Figure 2, Left). Existing approaches to counter early-stage entropy collapse rely on regularization techniques that merely slow the decay (Wang et al., 2025a; Yao et al., 2025; Zheng et al., 2025; Cheng et al., 2025). Our finding that one can actively increase policy entropy – while also improving validation performance – suggests a complementary strategy: using spurious-reward setups to more effectively preserve and modulate entropy. This highlights a promising direction for combining true and spurious rewards to better balance exploration and exploitation in RLVR.
@@ 5 Reward Misalignment: Who can Benefit from Random Rewards?
From empirical observations in this and prior work, we note two consistent patterns under random-reward training. First, in line with Shao et al. (2025), weaker models tend to improve less—and importantly, model strength is dataset-dependent: a model that performs well on an easier benchmark may struggle on a harder one. Second, as baseline accuracy increases (e.g., approaching 70%), training dynamics become noticeably smoother, whereas models starting around 50% accuracy exhibit substantially more oscillation. To explain when and why a model may improve under random rewards, we analyze the phenomenon through the lens of reward misalignment. As a warm-up, we introduce a simple probabilistic model that captures this mechanism in the binary outcome-reward (ORM) setting, converting the observed behavior into a tractable misalignment analysis.
@@
As reported by Shao et al. (2025), base Llama models reliably degrade during random-reward training across trials. Under the reward-misalignment perspective, stronger models should benefit more and are thus more likely to improve. We test this by evaluating a stronger distilled Llama variant, whose base and teacher models both exhibit contamination on MATH500. As shown in Figure 4 (Middle), using a rollout length of $8192$ tokens and matching all other hyperparameters to the Qwen-Math configuration, we observe improvements comparable to those in Figure 2. In contrast, the weaker and potentially contaminated Qwen-Math model (Figure 4, Left) fails to achieve similar gains. These results indicate that validation-set contamination does not account for the improvements under random rewards, nor is the effect specific to Qwen-Math. Figure 4 (Right) summarizes the percentage improvements across the model results in Figure 1 (Left) and Figure 4 (Left and Middle).
@@ Appendix A Related Works
We provide broader context on how prior work in reinforcement learning for classical settings has leveraged spurious rewards to facilitate training. Spurious reward signals are closely linked to challenges in generalization (Zhang et al., 2018; Koch et al., 2021; Langosco Di Langosco et al., 2022). While these works illustrate deliberate uses of such signals, spurious rewards may also arise unintentionally, leading to reward misspecification and reward hacking (Pan et al., 2022); similar reward hacking has been documented in Tien et al. (2023). A second relevant thread traces back to potential-based reward shaping (PBRS) (Ng et al., 1999), which introduces additional or misaligned rewards in principled ways to preserve the optimality of desired behaviors. More recently, Random Network Distillation (RND) (Burda et al., 2019) emerged as a leading exploration mechanism, with subsequent extensions such as Ma et al. (2025). In this same context, numerous other works propose reward signals (often spurious with respect to the true task objective) that encourage an agent to explore the state space in ways that eventually uncover genuine rewards (Pathak et al., 2017; Zhang et al., 2021; Wang et al., 2023; Li et al., 2024a). Spurious rewards have thus played a substantial role in improving exploration. One prominent theoretical foundation is posterior sampling (Russo and Van Roy, 2014), which motivates exploration through uncertainty and has been generalized to broader RL settings (Chen et al., 2024b; Xu et al., 2025).
@@
We now turn to recent works that study spurious rewards in RLVR. Although these works report broadly similar empirical phenomena, their experimental configurations differ in important ways. In Shao et al. (2025), the prompt omits the standard Qwen-style instruction to place the final answer in a box. As they note, Qwen-Math is highly sensitive to prompt formatting, and such differences can substantially shift baseline performance. In contrast, our experiments follow the default Qwen prompt used in verl (Sheng et al., 2025), which explicitly instructs the model to place the final answer in a boxed expression—mirroring the RLVR verifier in verl, which extracts the boxed answer for scoring and reward assignment.
Apart from this prompt choice, we match all other hyperparameters in Shao et al. (2025). Oertell et al. (2025), however, adopt a markedly different configuration: (i) a rollout-length cap of 1024 tokens (well below the 4096-token context window of Qwen-Math), (ii) a different training dataset (MATH (Hendrycks et al., 2021) instead of DeepScaleR (Luo et al., 2025)), (iii) a significantly smaller learning rate ($1\times 10^{-7}$ versus $5\times 10^{-7}$ in Shao et al. (2025)), and (iv) a reduced batch size (64 versus 128). The smaller learning rate changes the effective update magnitude, and the smaller batch size yields noisier estimates of the stochastic reward distribution. Given these differences, the empirical results reported across prior works are not directly comparable.
@@
We further comment on potential contamination in the Qwen2.5-Math models. As reported by Wu et al. (2025), contamination in Qwen-Math has been observed only on validation sets (e.g., MATH500). Beyond these findings, the official Qwen2.5-Math technical report (Yang et al., 2024, Table 1) also acknowledges possible contamination arising from the close similarity between the training and validation sets of the MATH dataset (Hendrycks et al., 2021). Their training corpus comprises two components: (i) CoT data synthesis, which includes GSM8K, MATH, and NuminaMath (Yang et al., 2024, §3.1.1), and (ii) a tool-integrated reasoning dataset containing GSM8K, MATH, CollegeMath, NuminaMath, MuggleMath, and DotaMath (Yang et al., 2024, §3.1.2). In contrast, our experiments employ the DeepScaleR training set, which consists exclusively of selected questions from AMC, AIME, Omni-Math, and Still (Luo et al., 2025). None of these datasets appear in the training sources listed for Qwen2.5-Math. Therefore, we believe that our training data does not overlap with the datasets used to train Qwen2.5-Math and is thus not contaminated.
@@
Agarwal et al. (2025) demonstrate that token-level entropy minimization can substantially improve LLM reasoning without verifiable feedback, arguing that reduced entropy increases model confidence and reveals latent reasoning capability. This mechanism parallels clipped training under random rewards, where updates primarily modulate entropy rather than exploit informative rewards. However, we show that entropy minimization alone may drive the policy toward low-entropy yet suboptimal solutions; hence, entropy should be viewed as a stabilizing regularizer rather than a replacement for genuine RLVR signals.
Related work explores entropy through the lens of self-confidence. In particular, Prabhudesai et al. (2025) use low-entropy rollouts as implicit rewards, achieving gains across diverse benchmarks, while Gao et al. (2025) show that even a single unlabeled example can improve reasoning via entropy reduction. Methods such as EMPO (Zhang et al., 2025b) and Zhao et al. (2025b) similarly enhance performance in unsupervised settings by amplifying model confidence. van Niekerk et al. (2025) further construct preference datasets from confidence scores, achieving RLHF-level improvements without human feedback. In this context, Cui et al. (2025) propose a simple but influential empirical relationship between policy entropy $\mathcal{H}$ and model performance $R$, fit across extensive experiments:
@@
To mitigate early-stage entropy collapse, several works propose alternative strategies. Shen (2025) analyze why entropy regularization suffers from limited benefit in RLVR training for LLMs by attributing it to the vast response space and the sparsity of optimal outputs, and then introduce an adaptive entropy-control method using a clamped entropy bonus with automatically tuned coefficients. Song et al. (2025) show that ORM yields induces sharp reductions in output entropy and diversity (as shown by lower pass@n scores), and propose outcome-level entropy bonuses to counteract it. Prior works (Wang et al., 2025a; Yao et al., 2025; Zheng et al., 2025; Cheng et al., 2025) also develop additional techniques for controlling entropy during RLVR training.
@@
Proximal policy optimization (PPO) (Schulman et al., 2017) has become standard for reward-based policy updates in LLM training and remains a core component of RLHF. However, since PPO requires loading and maintaining four separate models during training, it is computationally and memory intensive. This has motivated the development of lighter-weight and adapted policy-gradient updates (Li et al., 2024b; Ahmadian et al., 2024; Shao et al., 2024; Guo et al., 2025). In parallel, advances in verifiable reward construction (Cobbe et al., 2021; Uesato et al., 2022; Zelikman et al., 2022; Singh et al., 2023; Hosseini et al., 2024; Lightman et al., 2024; Wang et al., 2024; Luo et al., 2024; Setlur et al., 2025; Zhang et al., 2025c) have enabled reinforcement learning to substantially improve LLM reasoning, particularly in mathematical problem solving. Beyond training algorithms, recent work also explores post-processing and collaborative strategies to strengthen reasoning performance. Kay et al. (2025) and Zhao et al. (2025a) propose consensus-based and answer-aggregation methods within multi-model frameworks. Chen et al. (2025a) introduce a self-questioning paradigm for iterative refinement, while Park et al. (2025) develop an online multi-agent collaborative reinforcement learning framework.